OpenStudy (anonymous):

if sect= -5 and terminal point for t is in quadratic 2 find sin^2t+cos^2t

zepdrix (zepdrix):

Do they want us to work this out the long way? Or can we simply make use of our most basic trig identity?\[\Large\bf\sf \sin^2t+\cos^2t\quad=\quad 1\]

OpenStudy (anonymous):

it just says to find sin^2t+cos^2t

zepdrix (zepdrix):

Ok fine we'll do it the long way ^^

OpenStudy (anonymous):

which ever way makes more sense so I can understand it :)

zepdrix (zepdrix):

We can relate secant back to cosine using our identity,\[\Large\bf\sf \sec t\quad=\quad \frac{1}{\cos t}\] So if,\[\Large\bf\sf \sec t\quad=\quad -5\]It implies,\[\Large\bf\sf \cos t\quad=\quad -\frac{1}{5}\]That make sense so far? :o

OpenStudy (anonymous):

ya

zepdrix (zepdrix):

We're going to use this information to draw a triangle in the `second quadrant`. Because that's where they told us we're located.

zepdrix (zepdrix):